OpenStudy (anonymous):

Could someone help please a cellular company offers a contract for which the cost C, in dollars of t minutes of telephoning is given by C=0.25(t - 300) + 80.95, where it is assumed that t > or = to 300 minutes. what times will keep costs between $125.45 and $151.70? For the cost to be between $125.45 and $151.70, the telephoning time must be between ----minutes and ----minutes

OpenStudy (anonymous):

c=0.25(t-300)+ 80.95 when c= 125.45, 125.45=0.25(t-300)+80.95 44.5=0.25(t-300) 0.25t=44.5+75 t=478 min 151.70=0.25(t-300)+80.95 70.75+75= 0.25t t=583

OpenStudy (anonymous):

thanks c0c0ple the final answer is 478 minutes and 583 minutes correct

OpenStudy (anonymous):

ur welcome ^^